OpenStudy (anonymous):

Proof: cos² (a + b) + cos² (a - b) = 1 + cos 2a * cos 2b

OpenStudy (anonymous):

Resolve LHS using double angle formula of cosine and then use sum of angle formula of cosine to simplify.

OpenStudy (anonymous):

As english isn't my native language, with double angle formula you mean LHS = cos²(2a) = 1 + cos2a? But then there's no "b" left in the equation, so I'm probably going wrong somewhere at working out the formula :/

OpenStudy (anonymous):

see http://mathworld.wolfram.com/Double-AngleFormulas.html convert (cos(a))^2 into cos(2a).

OpenStudy (anonymous):

I'm sorry , i don't really get what you mean. This is all really new material to me and let's say I can't really rely on my teacher to explain it properly.. So basically the left side: cos²(a+b) + cos²(a-b) can be resolved into = cos² a = (1 + cos 2a)/2 right side: worked out by the double angle formulas? That does leave me with a "b" that is not in my left side of the equation...

OpenStudy (anonymous):

see the trouble is that on LHS the form is {cos(sum)}^2 + {cos(difference)}^2 while on RHS it is something to do with cos(double a) and cos(double b).So you need to make both sides in same form to compare.so convert cos(x)^2 terms of LHS into cos(2x) form.Here x will be a sum or difference.so thereafter use cos(m+n)=... formula.

OpenStudy (anonymous):

that gives me: 1 + cos[2(a+b)] + 1 + cos[2(a-b)]. = 1 + cos 2a * cos 2b - sin 2a * sin 2b + 1 + cos2a * cos2b + sin2b * sin2a = 1 + cos2a * cos2b + 1 + cos2a * cos2b = 2 + cos2a * cos 2b + cos2a * cos2b = 2 + 2(cos2a*cos2b) = 1 + cos2a * cos2b Is this correct?

OpenStudy (anonymous):

Actually, I think I forgot the division by 2 in the formula, adding that gives me the reason to divide by 2 at the 2 + 2(cos2a * cos2b) stage, I suppose?

OpenStudy (anonymous):

yep there should be a 1/2 in the first step itself.rest is correct..

OpenStudy (anonymous):

thanks